OpenStudy (anonymous):

Julia owns a hardware store. For every increase of 15cents in the price of a package of batteries, she estimates that sales decrease by 20 packages per day. The store normally sells 840 packages of batteries per day, at $6.00 per package. a) Determine an equation for the revenue, y, with x increase of 15cents b) What is the maximum daily revenue that Julia can expect from the battery sales? c) How many packages of batteries are sold when the revenue is at a maximum?

OpenStudy (anonymous):

for every \(x\) increase the price goes up by \(.15x\) and so the cost is \(6+.15x\) on the other hand for every \(x\) increase sales go down by 20 units from 840 so you sell \(840-20x\) units at a cost of \(6+.15x\) for a total revenue of \((840-20x)(6+.15x)=-3x^2+6 x+5040\) vertex (max) is when \(x=-\frac{b}{2a}=-\frac{6}{2\times -3}=1\) so max is at one 15 cent increase, sell for $6.15

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

you actually understood it?

OpenStudy (anonymous):

no i just made it up, could easily be wrong

OpenStudy (anonymous):

i was talking to EvaHeuves

OpenStudy (anonymous):

I don't get where your getting −3x2+6x+5040 from ?

OpenStudy (anonymous):

Oh never mind..

OpenStudy (anonymous):

ok so let's start from the beginning

OpenStudy (anonymous):

yess, so we got the whole y=(840-20x)(6+0.15x)

OpenStudy (anonymous):

you understood that part? @EvaHeuves

OpenStudy (anonymous):

yeah we did it in class together

OpenStudy (anonymous):

ok so then you want to foil that out

OpenStudy (anonymous):

Yes because didn't you try plugging a number into the equation and then you got your max value

OpenStudy (anonymous):

what's wrong with foil? easy mnemonic @satellite73

OpenStudy (anonymous):

y=5040+126x-120x-3x^2

OpenStudy (anonymous):

when you multiply out you get a parabola with negative leading coefficient, so it opens down and therefore has a maximum at the vertex. first coordinate of the vertex of any quadratic is always at \(x=-\frac{b}{2a}\) in this case it happened to be 1

OpenStudy (anonymous):

ok perfect @JeremySeed , now write that in standard form and then you can use the formula that @satellite73 used to find the vertex

OpenStudy (anonymous):

or y=-3x^2+6x+5040

OpenStudy (anonymous):

@EvaHeuves are yo ufollowing?

OpenStudy (anonymous):

ok there is nothing wrong with foil.

OpenStudy (anonymous):

Yes I am

OpenStudy (anonymous):

y=-3(x^2+6x)+5040 y=-3(x^2+6x+9-9)+5040 y=-3(x^2+6x+9)+5,067 y=-3(x+3)^2+5067

OpenStudy (anonymous):

good job @JeremySeed so now in this equation, what do x and y represent?

OpenStudy (anonymous):

somethings wrong cause that means the vertex is (-3,5067) meaning that when x is -3 the max value is reached

OpenStudy (anonymous):

you didn't factor the -3 out of the 6x

OpenStudy (anonymous):

ahhhhhh I get fully understand now

OpenStudy (anonymous):

how do you factor the -3 out of the 6x?

OpenStudy (anonymous):

it would be y=-3(x^2+2x)-5040

OpenStudy (anonymous):

oh daaaaaaaaaaaaaaaaaaaaaaaamnnnnnnnnnnnnnnnnnnnnnnnnnnnit

OpenStudy (anonymous):

k well we got the right idea on how to do it thats all that matters :)

OpenStudy (anonymous):

Yes, good job :)

OpenStudy (anonymous):

good job to you guys :) y'all did the math

OpenStudy (anonymous):

Thank you, and yes FOILING does help

OpenStudy (anonymous):

wow i didn't know this was so hard. forget this complete the square nonsense when you want the vertex (although it is very handy when you want to solve a quadratic) use the simple \(-\frac{b}{2a}\) as the \(x\) coordinate of the vertex, and for the \(y\) coordinate, substitute \(-\frac{b}{2a}\) for \(x\)

OpenStudy (anonymous):

we are in grade 10 sorry bud havent learned how to do that yet

OpenStudy (anonymous):

we have to do it the long way :/

OpenStudy (anonymous):

?? do it with \(ax^2+bx+c\) and see that you always get the same thing, namely \(-\frac{b}{2a}\)

OpenStudy (anonymous):

Yeah and you lost me.

OpenStudy (anonymous):

@satellite73 they have a test tomorrow, and it's best they do what their teacher taught them - not entirely sure if they get credit for work or not but either way go with what you already know guys for now

OpenStudy (anonymous):

@cerezas thank you so much for your help, much appreciated.

OpenStudy (anonymous):

then if you want to impress your teacher (the one that foisted "foil" on you) you can always write the answer as \(y=a(x-\frac{b}{2a})^2+\text{whatever}\) and you get the "whatever" by substitution

OpenStudy (anonymous):

id rather not try to impress my teacher, but forget how to do it and screw up, on something worth 10% of my final mark

OpenStudy (anonymous):

@EvaHeuves anytime, honestly and @satellite73 i agree it's a useful equation, now's just not the time

OpenStudy (anonymous):

@cerezas do you realize you have been helping me for the past 4 hours? You have no idea how much I appreciate all your help and how thankful I am for people like you :) you've actually been the biggest help i could ask for so thank you so much <3 :D

OpenStudy (anonymous):

couldnt*

OpenStudy (anonymous):

wait no could* :p

OpenStudy (anonymous):

@JeremySeed don't worry about it, i love doing this. i'm here whenever you need me seriously, i love doing this tell me how it goes tomorrow!

OpenStudy (anonymous):

same goes for you @EvaHeuves

OpenStudy (anonymous):

it's always great to help people who want to understand how to work through a problem and not necessarily just to get the answer <3

OpenStudy (anonymous):

yeah ill be sure to let you know! im sure it will be a lot better than it would have been if you never helped me :D

OpenStudy (anonymous):

awesome! like i said, i'm here all the time, and i can help in other subjects if you need it :D

OpenStudy (anonymous):

:) :) Did you like come from Heaven ? Like omg, I thought I was going to die but now I fully understand all the questions that you and @JeremySeed answered together. Thanks again :) <3

OpenStudy (anonymous):

glad to be of service! :)

OpenStudy (anonymous):

waaaait do you think youre up for answering one more? it was one i already asked and when the people answered i wasnt really clear on it..

OpenStudy (anonymous):

yeah sure just post it :)